Project

General

Profile

11-Oct-2018

Agenda:
1. General status updates/Schedule review
Rick -- DTC/CFO work in Rick’s office but not at test stand. Thinking was cooling is insufficient. 65C in office, 90C in bay. Added adhoc fans and got the bay ones down 60C. Then realized computer mother boards are different.. Checked power.. 12V rail provides all power. Tried external 12V power and it seemed to improve.. Then closed and nothing worked. Then swapped DTC and started working again. Made some changes to chants… maybe is reliable now.

Rick and John worked on cooling solultions.. Tying heat sink fan. Power two pin on board.

Rick is looking at reading other temperatures to indicate the problems.

Eric -- Adapted and integrated CFO run plan compiler into PCI linux kernel repo. Made a couple executables. Working with Tomo, exchanging emails, trying to get the analysis suite running.. Instrumenting profiler on art.

Bertrand -- Two modules take most of the time: first one (making data products) and clustering.. So asked Tomo to profile the two modules.

Jose -- Timing constraints seem to be more stable now. Attempting to get a DCS packet write/read to work. Getting weird data.

Greg -- Watching Rick suffer. Wants to hammer DTC/CFO link testing. Eric made it easy to access drivers (Eric says the test stand is intended to run autonomously,.. We could get back to that eventually).

IPR questions:
Recommendation pair-wise testing..

Ryan -- added ability to iterate on Configure.

We are OK with green lighting Dave Huffman for production on slow controls rack box.

2. Timing distribution summer work updates
3. online Trigger Processing (including through DTC) progress/plans
4. Status of tests with detector subsystems
5. Rack protection & slow controls progress/plans (dedicated DCS node at FCC)
Glenn -- beaglebone seems to be stable.